# Optimistic Safety for Linearly-Constrained Online Convex Optimization

### v1

@@ A.2 High probability events
First, we study the scaling that is required to take any $x\in\mathcal{Y}_{j}^{o}$ into $\mathcal{Y}_{j}^{p}$. It holds (by definition) that
$\hat{a}_{j,i}^{\top}x-\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}\leq b_{i}\quad\Longleftrightarrow\quad\hat{a}_{j,i}^{\top}x\leq b_{i}+\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}.$
Then, choosing $\alpha_{i}=\frac{b_{i}}{b_{i}+2\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}}$ such that $b_{i}\alpha_{i}=b_{i}-2\alpha_{i}\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}$ and $\alpha_{i}\in[0,1]$, we get that
$\hat{a}_{j,i}^{\top}(\alpha_{i}x)\leq\alpha_{i}\left(b_{i}+\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}\right)=\alpha_{i}b_{i}+\alpha_{i}\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}=b_{i}-\bar{\beta}_{j}\|(\alpha_{i}x)\|_{\bar{V}_{j}^{-1}}.$
This means that $\alpha_{i}x$ satisfies the $i$th constraint of the pessimistic set. Therefore, we can ensure that it holds for all constraints by choosing
$\bar{\alpha}:=\min_{i\in[n]}\left(\frac{b_{i}}{b_{i}+2\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}}\right)=\frac{b_{\mathrm{min}}}{b_{\mathrm{min}}+2\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}},$
where we use the fact that $\frac{b_{i}}{b_{i}+2\bar{\beta}_{j}\|x\|_{\bar{V}_{j}^{-1}}}$ is monotone in $b_{i}$. Since $\bar{\alpha}\in[0,1]$, $\mathbf{0}\in\mathcal{X}$, and $\mathcal{X}$ is convex, it holds that $\bar{\alpha}x\in\mathcal{Y}^{p}_{j}$. Because $\tilde{x}_{t}$ is in $\mathcal{Y}_{j_{t}}^{o}$, we can apply $x\leftrightarrow\tilde{x}_{t}$ to get that $\bar{\alpha}\tilde{x}_{t}$ is in $\mathcal{Y}^{p}_{j_{t}}$. Since $\gamma_{t}$ is the largest such scaling, $\bar{\alpha}$ is a lower bound on $\gamma_{t}$, i.e.
$\gamma_{t}=\max\{\alpha\in[0,1]:\alpha\tilde{x}_{t}\in\mathcal{Y}_{j_{t}}^{p}\}\geq\frac{b_{\mathrm{min}}}{b_{\mathrm{min}}+2\bar{\beta}_{j_{t}}\|\tilde{x}_{t}\|_{\bar{V}_{j_{t}}^{-1}}}.$ (4)
Since, $x_{t}=\gamma_{t}\tilde{x}_{t}$ and $\gamma_{t}\geq 0$, it follows that $\gamma_{t}\|\tilde{x}_{t}\|_{\bar{V}_{j_{t}}^{-1}}=\|\gamma_{t}\tilde{x}_{t}\|_{\bar{V}_{j_{t}}^{-1}}=\|x_{t}\|_{\bar{V}_{j_{t}}^{-1}}$. Therefore, we can rearrange (4) to get that
$\gamma_{t}\geq 1-2\frac{1}{b_{\mathrm{min}}}\bar{\beta}_{j_{t}}\|x_{t}\|_{\bar{V}_{j_{t}}^{-1}},$
@@
With this, we then restate and prove Lemma 4.2.
Lemma A.3 (Duplicate of Lemma 4.2).
@@
$\sum_{t=1}^{T}\left(f_{t}(\tilde{x}_{t})-f_{t}(x_{t})\right)$ $\leq G\sum_{t=1}^{T}\|\tilde{x}_{t}-x_{t}\|$
$\leq G\sum_{t=1}^{T}\|\tilde{x}_{t}-\gamma_{t}\tilde{x}_{t}\|$ (a)
$\leq DG\sum_{t=1}^{T}|1-\gamma_{t}|$
$\leq 2DG\frac{1}{b_{\mathrm{min}}}\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}\bar{\beta}_{j}\|x_{t}\|_{\bar{V}_{j}^{-1}}$ (b)
$\leq 2DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}\|x_{t}\|_{\bar{V}_{j}^{-1}}$ (c)
$\leq 4DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sum_{t=1}^{T}\|x_{t}\|_{V_{t}^{-1}}$ (d)
$\leq 4DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sqrt{T\sum_{t=1}^{T}\|x_{t}\|_{V_{t}^{-1}}^{2}}$ (e)
$\leq 4DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sqrt{2dT\log\left(1+\frac{T}{\lambda d}\right)}$ (f)
$\leq 4DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sqrt{3dT\log\left(T\right)},$ (g)
where (a) uses the fact that $x_{t}=\gamma_{t}\tilde{x}_{t}$, (b) follows from Lemma A.2, (c) uses the fact that $\beta_{T}\geq\beta_{t}$ for all $t\in[T]$, (d) uses the fact that for pd matrices $B\preceq C$ it holds that $\|z\|_{B^{-1}}\leq\frac{\det(C)}{\det(B)}\|z\|_{C^{-1}}$ for any $z$ (see e.g. Lemma 12 in Abbasi-Yadkori et al. (2011) or Lemma 27 in Cohen et al. (2019)), (e) is Cauchy-Schwarz, (f) follows from the elliptic potential lemma (Lemma A.6) and (g) is from the fact that $2d\log\left(1+\frac{T}{\lambda d}\right)\leq 2d\log\left(1+T\right)\leq 3d\log\left(T\right)$ when $T\geq 3$ and $\lambda=\max(1,D^{2})$. ∎
A.4 Proof of Lemma 4.3 (Optimistic Regret)
We restate and prove Lemma 4.3 in the following.
Lemma A.4 (Duplicate of Lemma 4.3).
@@
Let $x^{\star}_{j}$ be an optimal action within the optimistic set in phase $j$, i.e. $x^{\star}_{j}\in\operatorname*{arg\,min}_{x\in\mathcal{Y}_{j}^{o}}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x)$. Then, conditioned on $\mathcal{E}_{\mathrm{conf}}$, it holds that $x^{\star}\in\mathcal{Y}\subseteq\mathcal{Y}_{j}^{o}$ and therefore the optimal cost in each phase lower bounds the cost of the optimal action in each phase, i.e.
$\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star}_{j})\leq\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star}).$ (6)
Then, to apply the regret guarantee of $\mathcal{A}$, we need that the expectation over the distribution chosen by $\mathcal{A}$ is equal to the expectation conditioned on the randomness in previous rounds. Indeed, the distribution chosen by $\mathcal{A}$ in round $t$ (denoted $p_{t}$) is fully determined by the randomness in previous rounds (and $f_{t}$ is fixed) so it holds that $\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]=\mathbb{E}_{\tilde{x}_{t}\sim p_{t}}[f_{t}(\tilde{x}_{t})]$, and therefore
$\sum_{t=t_{j}}^{t_{j+1}-1}\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]-\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star}_{j})=\sum_{t=t_{j}}^{t_{j+1}-1}\mathbb{E}_{\tilde{x}_{t}\sim p_{t}}[f_{t}(\tilde{x}_{t})]-\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star}_{j})\leq C_{\mathcal{A}}\sqrt{t_{j+1}-t_{j}}$ (7)
for all $j\in[N]$. Also, note that the duration of each phase ($t_{j+1}-t_{j}$) depends on the actions chosen in that phase. Nonetheless, $\mathcal{A}$ is specified to handle a setting that can be terminated at any time and therefore the regret guarantees hold even when the duration of the phase depends on the actions taken.
Then, conditioned on $\mathcal{E}_{\mathrm{conf}}$, it holds that
$\sum_{t=1}^{T}\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]-\sum_{t=1}^{T}f_{t}(x^{\star})$ $=\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]-\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star})$
$\leq\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]-\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star}_{j})$ (a)
@@
where (a) uses (6), (b) uses (7), (c) is from the specification that $\mathcal{A}$ is polylogarithmic in $t_{j+1}-t_{j}$ and therefore (with abuse of notation) pull it out of the summation to mean that it is polylogarithmic in $T$ as $t_{j+1}-t_{j}\leq T$, (d) is Cauchy-Schwarz and (e) uses $N\leq 4d\log(T)$ from Lemma A.7.
Then, also conditioning on $\mathcal{E}_{\mathrm{azuma}}$, it holds
$\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})-\sum_{t=1}^{T}f_{t}(x^{\star})$ $=\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})-\sum_{t=1}^{T}\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]+\sum_{t=1}^{T}\mathbb{E}_{t}[f_{t}(\tilde{x}_{t})]-\sum_{t=1}^{T}f_{t}(x^{\star})$
$=2DG\sqrt{2T\log(1/\delta)}+C_{\mathcal{A}}\sqrt{4d\log(T)T},$
@@
A.5 Completing the proof
We restate and prove Theorem 4.1 in the following.
Theorem A.5 (Duplicate of Theorem 4.1).
Fix some $\delta\in(0,1/2)$. Then, with probability at least $1-2\delta$, the regret of OSOCO (Algorithm 1) satisfies
$R_{T}\leq 4DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sqrt{3dT\log\left(T\right)}+C_{\mathcal{A}}\sqrt{4dT\log(T)}+2DG\sqrt{2T\log(1/\delta)}$

### v3

@@ B.3 Term I (Cost of Safety)
Throughout the proof, we we take $j=j_{t}$ to simplify notation. Then to establish the proof, we first show that $\alpha\tilde{x}_{t}\in\mathcal{Y}_{j}^{p}$ where
$\alpha:=\frac{b_{\mathrm{min}}}{b_{\mathrm{min}}+2\bar{\beta}_{j}\|\tilde{x}_{t}\|_{\bar{V}_{j}^{-1}}}$
and then complete the proof by showing that this implies the statement of the lemma. Each of these steps are given in the following sections.
Proof that $\alpha\tilde{x}_{t}\in\mathcal{Y}_{j}^{p}$:
It holds that $\alpha\tilde{x}_{t}\in\mathcal{Y}_{j}^{p}$ if both (a) $\alpha\tilde{x}_{t}\in\mathcal{X}$ and (b) $\hat{a}_{j,i}^{\top}(\alpha\tilde{x}_{t})+\bar{\beta}_{j}\|\alpha\tilde{x}_{t}\|_{\bar{V}_{j}^{-1}}\leq b_{i}$ for all $i\in[n]$. Point (a) holds because $\alpha\in[0,1]$, $\tilde{x}_{t}\in\mathcal{X}$, $\mathbf{0}\in\mathcal{X}$ and $\mathcal{X}$ is convex, and therefore
$\alpha\tilde{x}_{t}=\alpha\tilde{x}_{t}+(1-\alpha)\mathbf{0}\in\mathcal{X}.$
Point (b) holds given that for all $i\in[n]$,
$\hat{a}_{j,i}^{\top}(\alpha\tilde{x}_{t})+\bar{\beta}_{j}\|\alpha\tilde{x}_{t}\|_{\bar{V}_{j}^{-1}}$ $=\alpha(\hat{a}_{j,i}^{\top}\tilde{x}_{t}+\bar{\beta}_{j}\|\tilde{x}_{t}\|_{\bar{V}_{j}^{-1}})$
$=\alpha(\hat{a}_{j,i}^{\top}\tilde{x}_{t}-\bar{\beta}_{j}\|\tilde{x}_{t}\|_{\bar{V}_{j}^{-1}}+2\bar{\beta}_{j}\|\tilde{x}_{t}\|_{\bar{V}_{j}^{-1}})$
$\leq\alpha(b_{i}+2\bar{\beta}_{j}\|\tilde{x}_{t}\|_{\bar{V}_{j}^{-1}})$ (a)
$=\frac{b_{\mathrm{min}}}{b_{\mathrm{min}}+2\bar{\beta}_{j}\|\tilde{x}_{t}\|_{\bar{V}_{j}^{-1}}}(b_{i}+2\bar{\beta}_{j}\|\tilde{x}_{t}\|_{\bar{V}_{j}^{-1}})$ (b)
$\leq\frac{b_{i}}{b_{i}+2\bar{\beta}_{j}\|\tilde{x}_{t}\|_{\bar{V}_{j}^{-1}}}(b_{i}+2\bar{\beta}_{j}\|\tilde{x}_{t}\|_{\bar{V}_{j}^{-1}})$ (c)
$=b_{i}$
where (a) uses the fact that $\tilde{x}_{t}\in\mathcal{Y}_{j}^{o}$, (b) uses the definition of $\alpha$, and (c) uses the fact that $\frac{b_{\mathrm{min}}}{b_{\mathrm{min}}+2\bar{\beta}_{j}\|\tilde{x}_{t}\|_{\bar{V}_{j}^{-1}}}$ is increasing in $b_{\mathrm{min}}$ and $b_{\mathrm{min}}\leq b_{i}$.
Completing the proof:
From the definition of $\gamma_{t}$, it holds that
$\gamma_{t}=\max\{\alpha\in[0,1]:\alpha\tilde{x}_{t}\in\mathcal{Y}_{j}^{p}\}\geq\frac{b_{\mathrm{min}}}{b_{\mathrm{min}}+2\bar{\beta}_{j}\|\tilde{x}_{t}\|_{\bar{V}_{j}^{-1}}}.$ (5)
Then, since $x_{t}=\gamma_{t}\tilde{x}_{t}$ and $\gamma_{t}\geq 0$, it also holds that
$\gamma_{t}\|\tilde{x}_{t}\|_{\bar{V}_{j}^{-1}}=\|\gamma_{t}\tilde{x}_{t}\|_{\bar{V}_{j}^{-1}}=\|x_{t}\|_{\bar{V}_{j}^{-1}}.$
Therefore, we can rearrange (5) to get that
$\gamma_{t}\geq 1-2\frac{1}{b_{\mathrm{min}}}\bar{\beta}_{j}\|x_{t}\|_{\bar{V}_{j}^{-1}},$
@@
With this, we then give the bound on Term I.
Lemma 5 (Cost of Safety).
@@
$\sum_{t=1}^{T}\left(f_{t}(x_{t})-f_{t}(\tilde{x}_{t})\right)$ $\leq G\sum_{t=1}^{T}\|\tilde{x}_{t}-x_{t}\|$
$=G\sum_{t=1}^{T}\|\tilde{x}_{t}-\gamma_{t}\tilde{x}_{t}\|$ (a)
$\leq DG\sum_{t=1}^{T}(1-\gamma_{t})$
$\leq 2DG\frac{1}{b_{\mathrm{min}}}\sum_{t=1}^{T}\bar{\beta}_{j_{t}}\|x_{t}\|_{\bar{V}_{j_{t}}^{-1}}$ (b)
$\leq 2DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sum_{t=1}^{T}\|x_{t}\|_{\bar{V}_{j_{t}}^{-1}}$ (c)
$\leq 2DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sum_{t=1}^{T}\frac{\det(V_{t})}{\det(\bar{V}_{j_{t}})}\|x_{t}\|_{V_{t}^{-1}}$ (d)
$\leq 4DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sum_{t=1}^{T}\|x_{t}\|_{V_{t}^{-1}}$ (e)
$\leq 4DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sqrt{T\sum_{t=1}^{T}\|x_{t}\|_{V_{t}^{-1}}^{2}}$ (f)
$\leq 4DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sqrt{2dT\log\left(1+\frac{T}{\lambda d}\right)}$ (g)
$\leq 4DG\frac{1}{b_{\mathrm{min}}}\beta_{T}\sqrt{3dT\log\left(T\right)},$ (h)
where each step is justified in the following:
  • (a)
    $x_{t}=\gamma_{t}\tilde{x}_{t}$ as specified in line 11 in the algorithm.
  • (b)
    Lemma 4.
  • (c)
    $\beta_{t}$ is increasing in $t$.
  • (d)
    For pd matrices $B\preceq C$ it holds that $\|z\|_{B^{-1}}\leq\frac{\det(C)}{\det(B)}\|z\|_{C^{-1}}$ for any $z$ (see e.g. Lemma 12 in Abbasi-Yadkori et al. (2011) or Lemma 27 in Cohen et al. (2019)).
  • (e)
    Follows from phase termination criteria (line 14) which ensures that $\det(V_{t})\leq 2\det(\bar{V}_{j_{t}})$ for all $t$.
  • (f)
    Cauchy-Schwarz.
  • (g)
    Elliptic potential lemma (Lemma 7).
  • (h)
    $2d\log\left(1+\frac{T}{\lambda d}\right)\leq 2d\log\left(1+T\right)\leq 3d\log\left(T\right)$ when $T\geq 3$ and $\lambda=\max(1,D^{2})$.
B.4 Term II (Optimistic Regret)
We prove a bound on Term II in the following.
Lemma 6 (Optimistic Regret).
@@ B.4 Term II (Optimistic Regret)
Let $x^{\star}_{j}$ be an optimal action within the optimistic set in phase $j$, i.e. $x^{\star}_{j}\in\operatorname*{arg\,min}_{x\in\mathcal{Y}_{j}^{o}}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x)$. Then, consider the decomposition,
$\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})-\sum_{t=1}^{T}f_{t}(x^{\star})=$ $\underbrace{\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})-\sum_{t=1}^{T}\mathbb{E}[f_{t}(\tilde{x}_{t})|\mathcal{H}_{t-1}]}_{\mathrm{Term\ I}}$
$+\underbrace{\sum_{t=1}^{T}\mathbb{E}[f_{t}(\tilde{x}_{t})|\mathcal{H}_{t-1}]-\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star}_{j})}_{\mathrm{Term\ II}}$
$+\underbrace{\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star}_{j})-\sum_{t=1}^{T}f_{t}(x^{\star}).}_{\mathrm{Term\ III}}$
We bound each of the terms in the following sections.
Term I:
Conditioning on $\mathcal{E}_{\mathrm{azuma}}$, it holds by definition that
$\mathrm{Term\ I}=\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})-\sum_{t=1}^{T}\mathbb{E}[f_{t}(\tilde{x}_{t})|\mathcal{H}_{t-1}]\leq 2DG\sqrt{2T\log(1/\delta)}.$ (7)
Term II:
We bound Term II as follows,
$\mathrm{Term\ II}$ $=\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}\mathbb{E}[f_{t}(\tilde{x}_{t})|\mathcal{H}_{t-1}]-\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star}_{j})$
$=\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}\mathbb{E}_{\tilde{x}_{t}\sim p_{t}}[f_{t}(\tilde{x}_{t})]-\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star}_{j})$ (a)
@@
where each step is justified in the following:
  • (a)
    Note that the the distribution from which $\tilde{x}_{t}$ is sampled (denoted by $p_{t}$) is fully determined by the randomness in rounds $[t-1]$ and therefore $p_{t}$ is the conditional distribution of $\tilde{x}_{t}$ given $\mathcal{H}_{t-1}$. Additionally, the adversary chooses $f_{t}$ simultaneous to the player choosing $x_{t}$ and therefore $f_{t}$ is determined by the randomness in rounds $[t-1]$. Therefore, the conditional distribution of $f_{t}(\tilde{x}_{t})$ given $\mathcal{H}_{t-1}$ is $p_{t}$, and thus $\mathbb{E}[f_{t}(\tilde{x}_{t})|\mathcal{H}_{t-1}]=\mathbb{E}_{\tilde{x}_{t}\sim p_{t}}[f_{t}(\tilde{x}_{t})]$.
  • (b)
    The regret bound of $\mathcal{A}$ as specified in Definition 1. Note that the duration of each phase ($t_{j+1}-t_{j}$) depends on the actions chosen in that phase. Nonetheless, $\mathcal{A}$ is specified to handle a setting that can be terminated at any time and therefore the regret guarantees hold even when the duration of the phase depends on the actions taken.
  • (c)
    The fact that $C_{\mathcal{A}}$ is polylogarithmic in $t_{j+1}-t_{j}$ and therefore (with abuse of notation) pull it out of the summation to mean that it is polylogarithmic in $T$ as $t_{j+1}-t_{j}\leq T$.
  • (d)
    Cauchy-Schwarz.
  • (e)
    $N\leq 4d\log(T)$ from Lemma 8.
Term III:
Conditioned on $\mathcal{E}_{\mathrm{conf}}$, it holds that
$\mathrm{Term\ III}$ $=\sum_{j=1}^{N}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x^{\star}_{j})-\sum_{t=1}^{T}f_{t}(x^{\star})$
$=\sum_{j=1}^{N}\min_{x\in\mathcal{Y}_{j}^{o}}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x)-\sum_{t=1}^{T}f_{t}(x^{\star})$
$\leq\sum_{j=1}^{N}\min_{x\in\mathcal{Y}}\sum_{t=t_{j}}^{t_{j+1}-1}f_{t}(x)-\sum_{t=1}^{T}f_{t}(x^{\star})$
$\leq\min_{x\in\mathcal{Y}}\sum_{t=1}^{T}f_{t}(x)-\sum_{t=1}^{T}f_{t}(x^{\star})=0,$
where the first inequality uses $\mathcal{Y}\subseteq\mathcal{Y}_{j}^{o}$ for all $j$ which follows from Theorem 2.
Completing the proof:
Putting everything together, we get, conditioned on $\mathcal{E}_{\mathrm{conf}}$ and $\mathcal{E}_{\mathrm{azuma}}$ that
$\sum_{t=1}^{T}f_{t}(\tilde{x}_{t})-\sum_{t=1}^{T}f_{t}(x^{\star})\leq C_{\mathcal{A}}\sqrt{4dT\log(T)}+DG\sqrt{2T\log(1/\delta)},$
@@
B.5 Completing the proof
Finally, we restate Theorem 1 in the following.
Theorem 3 (Restatement of Theorem 1).
Fix some $\delta\in(0,1/2)$. Then, with probability at least $1-2\delta$, the actions of OSOCO (Algorithm 1) satisfy,